OpenStudy (anonymous):

1/4p- 3/8p=4 solve for p

OpenStudy (anonymous):

p=-32

OpenStudy (anonymous):

EXPLANATION (1)/(4)*p-(3)/(8)*p=4 Multiply 1 by p to get p. (p)/(4)-(3)/(8)*p=4 Multiply -3 by p to get -3p. (p)/(4)-(3p)/(8)=4 Find the LCD (least common denominator) of (p)/(4)-(3p)/(8)+4. Least common denominator: 8 Multiply each term in the equation by 8 in order to remove all the denominators from the equation. (p)/(4)*8-(3p)/(8)*8=4*8 Cancel the common factor of 4 in the denominator of the first term (p)/(4) and the second term 8. (p)/(<X>4<x>)*<X>8<x>^(2)-(3p)/(8)*8=4*8 Reduce the expression by removing the common factor of 4 in the denominator of the first term (p)/(4) and the second term 8. p*2-(3p)/(8)*8=4*8 Cancel the common factor of 8 in the denominator of the first term -(3p)/(8) and the second term 8. p*2-(3p)/(<X>8<x>)*<X>8<x>=4*8 Reduce the expression by removing the common factor of 8 in the denominator of the first term -(3p)/(8) and the second term 8. p*2-3p*1=4*8 Multiply p by 2 to get 2p. 2p-3p*1=4*8 Multiply -3p by 1 to get -3p. 2p-3p=4*8 According to the distributive property, for any numbers a, b, and c, a(b+c)=ab+ac and (b+c)a=ba+ca. Here, p is a factor of both 2p and -3p. (2-3)p=4*8 To add integers with different signs, subtract their absolute values and give the result the same sign as the integer with the greater absolute value. In this example, subtract the absolute values of 2 and -3 and give the result the same sign as the integer with the greater absolute value. (-1)p=4*8 Remove the parentheses. -p=4*8 Multiply 4 by 8 to get 32. -p=32 Multiply each term in the equation by -1. -p*-1=32*-1 Multiply -p by -1 to get p. p=32*-1 Multiply 32 by -1 to get -32. p=-32

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcome! :D

OpenStudy (whpalmer4):

\[\frac{1}{4}p -\frac{3}{8}p=4\] I would multiply both sides by 8, which would eliminate the fractions (in general, the least common multiple of all of the denominators) \[8*\frac{1}{4}p -8*\frac{3}{8}p=8*4\]\[2p -3p=32\]\[-p=32\]\[p=-32\]